Two 110 kg lead spheres are suspended from 200 m -long massless cables. The tops of the cables have been carefully anchored exactly 1.0 m apart.

By how much is the distance between the centers of the spheres less than 1.0 meter?
Answer #1

Gravitational force

F=\frac{Gm^2}{r^2}=mg\tan \theta

Where

\tan \theta= \frac{Gm}{gr^2}

Distance change on each side L\sin \theta

Total distance s=2L \sin \theta

For very small angles \sin \theta = \tan \theta= \theta

S=2L \theta

S=2L \frac{Gm}{gr^2}=2(200)(\frac{6.67x10^{-11}*110}{9.8*(1)^2})

S= 2.99x10^{-7}m